### 27 Feb 2024 v1

# RISAR: RIS-assisted Human Activity Recognition with Commercial Wi-Fi Devices

Abstract
Human activity recognition (HAR) is integral to smart homes, security, and healthcare. Existing systems face challenges such as the absence of line-of-sight links, insufficient details regarding the sensing subject, and inefficiencies in noise reduction and feature extraction from channel state information. To address these issues, this study builds a reconfigurable intelligent surface (RIS)-assisted passive human activity recognition (RISAR) system, compatible with commercial Wi-Fi devices. RISAR employs a RIS to augment the orientation and spatial diversity of Wi-Fi signals, thereby facilitating improved detection capabilities. A new denoising and feature extraction technique, the high-dimensional factor model, based on random matrix theory, is proposed for noise elimination and salient temporal feature extraction. On this basis, a dual-stream spatial-temporal attention network model is developed for assigning variable weights to different characteristics and sequences, mimicking human cognitive processes in prioritizing essential information. Experimental analysis shows that RISAR significantly outperforms existing HAR systems in accuracy and efficiency, achieving an average accuracy of 97.26%. These findings not only highlight the adaptability of the RISAR system but also underscore its potential as a robust solution for activity recognition across complex environments.
Index Terms:
Reconfigurable intelligent surface, channel state information, human activity recognition, random matrix theory, neural network.
I Introduction
Human activity recognition (HAR) is a crucial component in the Internet of Things (IoT) realm, particularly in smart homes and offices [1]. It entails the classification of human activities, thereby enabling applications in eldercare, healthcare, and home security [2, 3, 4, 5]. HAR methodologies can be systematically categorized into three distinct domains, predicated on the sensing modalities employed: vision-based HAR [6, 7, 8, 9], wearable device-based HAR [10, 2, 11, 12], and wireless sensing-based HAR [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Vision-based HAR, which primarily employs high-resolution cameras, contends with impediments in non-line-of-sight (NLOS) scenarios and engenders privacy concerns [26]. Wearable device-based HAR, leveraging integrated sensors such as accelerometers, magnetometers, gyroscopes, barometers, and physiological sensors, faces the drawback of potential user inconvenience, particularly among the elderly and children, due to the necessity of continuous wear [26]. Wireless sensing-based HAR encompasses technologies like radar, radio frequency identification (RFID), and Wi-Fi signals. Radars, capable of detecting human motion without necessitating user-carried devices, offer resilience against variations in lighting conditions due to their broad bandwidth [27]. However, challenges in radar-based HAR systems include limitations in portability and elevated hardware costs. Conversely, RFID systems present a cost-effective alternative, requiring minimal hardware such as chip tags and antennas. Despite their economic advantages, RFID systems face robustness issues in intricate indoor environments, predominantly due to factors like multipath effect and temporal dynamics [28].
Recent advancements in Wi-Fi signal-based HAR techniques have elicited significant interest in the field. These methodologies leverage channel state information (CSI) and multi-antenna configurations to effectively distinguish between various human activities, as demonstrated in systems such as WiSee [15], RT-Fall [16], CARM [17], WiAct [20], and CeHAR [24]. For instance, RT-Fall [16] utilized both amplitude and phase data from CSI measurements to identify falls, while CARM [17] applied CSI-speed and CSI-activity models for recognizing activities. Furthermore, WiAct [20] explored the correlation between body movement and CSI amplitude to classify activities. Despite the passive detection capabilities and the ease of deployment characterizing these systems, they face notable challenges, particularly in NLOS scenarios and in lack of spatial diversity. These limitations underscore the necessity for ongoing research and development in HAR technologies to achieve robust performance across a wider array of environmental conditions.
In typical indoor environments, often comprised of multiple rooms linked to a single Wi-Fi access point (AP), Wi-Fi signals are frequently confronted with NLOS conditions. These scenarios are characterized by physical barriers and distance, which hinder both direct and reflected signal propagation paths between the transmitter and receiver. As a result, such environments experience elevated levels of noise, interference, and multipath effects [29]. Furthermore, traditional denoising and feature extraction methods, like principal component analysis (PCA), while effective for linear dimensionality reduction, may not adequately tackle the complexities associated with Wi-Fi signal data under these conditions [30]. Our empirical research indicates that PCA, in particular, delivers suboptimal performance in activity recognition within NLOS contexts. This finding underscores the imperative for investigating more sophisticated techniques to improve HAR performance in these demanding environments.
In this research, we present a novel device-free reconfigurable intelligent surface-assisted human activity recognition (RISAR) system that utilizes Wi-Fi signals, specifically designed for NLOS scenarios. The RISAR system comprises a pair of personal computers (PCs), with one functioning as the transmitter and the other as the receiver, in conjunction with RISs. Addressing the primary challenge, we integrate RIS technology into the HAR domain. Celebrated for its groundbreaking potential in the field of wireless communications, RIS enables the dynamic adaptation of propagation channels, thereby fostering an environment optimized for signal transmission [31]. This strategic enhancement facilitates the creation of multiple independent signal pathways, significantly boosting the accuracy of human activity recognition [32]. Moreover, the strategic deployment of RIS utilizes spatial diversity to effectively diminish the multipath effect, ensuring enhanced sensing performance [33, 34, 35]. To address the secondary challenge, we introduce a high-dimensional factor model for the extraction of activity features from the raw CSI measurements. Additionally, a specialized dual-stream spatial-temporal attention network (DS-STAN) architecture, expressly designed for the nuanced characteristics of CSI data, has been developed. This architecture is adept at processing and interpreting the complex patterns within CSI data, thereby enhancing our HAR system’s proficiency in the accurate identification and classification of human activities.
The key contributions of our research are delineated as follows:
  • $\bullet$
    A RISAR system is built, which capitalizes on Wi-Fi signals, especially for NLOS conditions. The incorporation of RIS not only introduces unique signal pathways but also enhances spatial diversity, offering a significant improvement in signal quality and system performance. To the best of our knowledge, this is the first work to employ RIS technology for human activity recognition with commodity Wi-Fi devices.
  • $\bullet$
    To enable the effective implementation of RISAR, a high-dimensional factor model approach is proposed for extracting activity features from CSI measurements. These extracted features are fed into the proposed DS-STAN model that integrates an attention mechanism to discriminate various human activities. This approach represents a sophisticated fusion of random matrix theory and deep learning techniques, offering enhanced accuracy in activity recognition.
  • $\bullet$
    The RISAR system, operational with commercial Wi-Fi devices under both LOS and NLOS conditions, has undergone extensive evaluation. The empirical results highlight the system’s remarkable accuracy, reaching an average of 97.26% in LOS scenarios and 90.83% in NLOS configurations. Such performance not only illustrates RISAR’s robustness but also its adaptability across various environmental settings, signifying a substantial leap forward in HAR technology.
The remainder of this paper is as follows: Section II encompasses a comprehensive literature review of pertinent existing works. Section III presents the system information. Section IV details our experimental results, and Section V concludes the paper.
II Related Works
In this section, we summarize the related work on human activity recognition and Wi-Fi signal-based activity recognition.
II-A Human Activity Recognition
Current endeavors in HAR predominantly encompass three distinct categories: vision-based [6, 7, 8, 9], wearable device-based [10, 2, 11, 12], and wireless sensing-based systems [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
Vision-based methodologies primarily leverage cameras to capture image sequences and videos for activity recognition [6]. For instance, Kim et al. [7] introduced a depth video-based HAR system, utilizing skeletal joint features to discern the daily activities of the elderly in indoor settings. Additionally, Franco et al. [8] amalgamated skeleton and RGB data stream analyses to extract features for HAR. Despite their effectiveness, these systems are susceptible to lighting conditions and obstructions, and they raise significant security and privacy concerns.
Wearable device-based HAR approaches utilize technologies such as smartphones [2], IoT devices [10], and sensors [11] for activity detection. Hassan et al. [2] developed a robust HAR system based on smartphone sensor data, capable of distinguishing transitional and non-transitional activities. Bhat et al. [10] employed a low-power IoT device for HAR, pioneering a framework for both online training and inference. Bianchi et al. [11] explored the integration of inertial measurement units (IMUs) with deep learning, where wearable sensors transmit data to a cloud service for activity monitoring. These methods, however, rely on active detection, necessitating the continuous wearing of the device.
In the realm of wireless sensing-based HAR, techniques utilize wireless signals for monitoring activities. Zhu et al. [13] proposed a distributed ultra-wideband (UWB) radar-based system, classifying motion events through spatio-temporal patterns of radar Doppler signatures. Oguntala et al. [14] introduced SmartWall, a system employing passive RFID tags for the recognition of sequential and concurrent activities. While these techniques obviate the need for wearing devices, they often require dedicated devices to be installed in the environment.
II-B Wi-Fi signal-based activity recognition
In the evolving landscape of Wi-Fi signal-based technologies, a diverse array of systems have been developed for HAR. Yet, these systems frequently necessitate specialized devices and are often inadequate in NLOS scenarios. For instance, the WiSee system, as demonstrated by Pu et al. [15], confirmed the efficacy of Wi-Fi signals in detecting human motion in NLOS and through-the-wall environments, albeit relying on USRP-N210 devices rather than standard commercial Wi-Fi hardware. Yousefi et al. [18] employed long short-term memory (LSTM) networks to encode temporal information during feature learning, setting a new benchmark in HAR using CSI measurements. Similarly, Chen et al. [19] innovatively utilized raw CSI signals in conjunction with an attention-based bi-directional LSTM to ascertain action classes. Feng et al. [21] proposed a three-phase system for multiple human activity recognition, focusing on identifying the start and end points of activities in multi-subject scenarios. Xiao et al. [22] introduced a deep learning-based HAR approach that capitalizes on Wi-Fi signals, transforming segmentation tasks into classification challenges via a convolutional neural network (CNN) algorithm. Guo et al. [23] developed an encoder-decoder framework aimed at mitigating accuracy disparities among individuals in HAR applications. Lu et al. [24] in their CeHAR system, advocated for the amalgamation of CSI amplitude and phase features to enhance action detection efficiency. Lastly, Shalaby et al. [25] presented a deep learning model integrating CNNs, gated recurrent units (GRUs), and LSTMs, designed to extract high-dimensional and time sequence features.
Despite their advantages in passive detection, ease of deployment, and cost-effectiveness, these systems generally do not consider HAR in NLOS environments. This is particularly pertinent in many indoor settings, where multiple rooms often connect to a single wireless signal AP, especially in residential environments. In such scenarios, issues like signal attenuation and multipath fading can significantly impair the accuracy of activity recognition.
III System Framework
The architecture of the RISAR system, as presented in Fig. 1, is comprised of several integral stages, namely: CSI data collection, data preprocessing, and human activity classification.
Figure 1: The framework of the proposed RISAR system.
III-A RISAR System
In contrast to traditional radio-frequency (RF) sensing methods, which are often constrained by their radio environments and limited in the number of transmission channels for multi-dimensional human activity information, our research introduces the RISAR system. This system is specifically designed for human activity recognition using RIS in NLOS environments. The data collection for activity recognition was conducted in three distinct environments, as depicted in Fig. 2. The RISAR system is comprised of two PCs, each equipped with an Intel Wi-Fi Link 5300 network interface card (NIC), functioning as the transmitter (Tx) and receiver (Rx). In environment A, the setup includes a PC with a horn antenna as the Tx and a PC with three antennas as the Rx. In environments B and C, both Tx and Rx are configured with three antennas each. When a human activity occurs within the transmission and reception range, the RISAR system is capable of automatically detecting it.
(a) Environment A.
(b) Floor plans of two experimental environments.
Figure 2: Three types of CSI data collection environments.
Under NLOS conditions, as illustrated in Fig.2LABEL:sub@F2-a, the experimental setup involves a Tx PC outfitted with an Intel 5300 NIC and a horn antenna, positioned 1 meter high, directing signals towards the RIS located at a corner 5 meters away, thereby establishing a LOS link. The RIS, placed at the same 1-meter height, is optimized using the DaS algorithm [36] to adjust its phase matrix, maximizing signal power at the Rx location. Concurrently, the Rx PC, also equipped with an Intel 5300 NIC and an omnidirectional antenna, is stationed on a 1-meter-high tripod. The efficacy of the RIS, as detailed in a prior study by the authors [37], is validated, with the designed RIS operating at 5.8 GHz yielding an improvement of over 10 dBi. For LOS conditions, the experimental setup, as delineated in Fig.2LABEL:sub@F2-b, incorporates data collection in a meeting room within a $12\times 8.7\,\text{m}^{2}$ area and an office room covering an $8\times 8.7\,\text{m}^{2}$ space. The Intel 5300 NIC in our experiments was configured to operate in the 5 GHz band, utilizing the 161-st channel with a 20 MHz channel bandwidth. The RIS board consists of a uniform planar array of $10\times 16$ units, totaling 160, and operates at a central frequency of 5.8 GHz.
III-B Data Collection
CSI measurements are obtained from packets captured by the receiver. As delineated by Halperin et al. [38], each packet encompasses a CSI matrix of dimensions $N_{\text{Tx}}\times N_{\text{Rx}}\times N_{\text{sc}}$, where $N_{\text{Tx}}$ and $N_{\text{Rx}}$ signify the count of transmitting and receiving antennas, respectively, and $N_{\text{sc}}$ indicates the orthogonal frequency-division multiplexing (OFDM) subcarriers. Specifically, the Intel Wi-Fi Link 5300 NIC, which uses 56 subcarriers, exports CSI data for 30 subcarriers across three antennas, resulting in each packet incorporating $1\times 3\times 30$ CSI measurements for NLOS and $3\times 3\times 30$ for LOS scenarios.
In the proposed RISAR system, the CSI sampling rate is set at 1000 packets per second over a 5-second interval. We curate diverse datasets in environments labeled as RIS-corner (Environment A), Meeting room (Environment B), and Office (Environment C) to appraise the system’s capability in discerning a variety of daily activities. Each activity is executed 100 times by participants to amass a comprehensive dataset for analysis, as cataloged in Table I.
TABLE I: Typical categories of activities tested in the experiments.
Human activity Number of samples
RIS-corner Meeting room Office
bend 400 - 600
empty 400 - -
fall 400 - 600
lie down - 600 -
pick up - 600 -
run - - 600
sit down 400 600 600
stand - 600 600
stand up 400 600 -
walk - 600 600
wave 400 - 600
  • - denotes there is no sample in the dataset.
III-C Data Analysis and Preprocessing
Given that low-cost commercial Wi-Fi cards typically introduce diverse types of noise into CSI measurements [39], it necessitates a preprocessing phase for the CSI data. This preprocessing is vital to extract relevant features, substantially enhancing the effectiveness of algorithmic training and thus significantly contributing to the achievement of Wi-Fi sensing objectives.
The preprocessing phase of the RISAR system encompasses two critical components: the denoising and conversion of CSI data, followed by signal extraction. Initially, phase offset in CSI measurement values, including sampling time offset (STO) and sampling frequency offset (SFO), is addressed through a linear transformation method, as suggested in previous research [40]. This approach effectively mitigates shifts caused by STO/SFO, yielding more accurate phase values. Subsequently, our paper applies this method to ascertain precise phase values.
III-C1 Activity feature analysis
Raw CSI measurements primarily capture amplitude and phase changes over time but lack inherent frequency components. The fast Fourier transform (FFT) is a standard technique for converting CSI measurements from the time to the frequency domain, aiding in power spectral density (PSD) extraction but compromising time-domain information [41, 42]. In contrast, the short-time Fourier transform (STFT) and discrete wavelet transform (DWT) enable the capture of features in both time and frequency domains [16, 43, 44, 17]. Fig. 3 illustrates the STFT (spectrogram) of CSI for various activities, revealing that activities with significant movements, such as walking, display high energy in the spectrogram. This crucial time-frequency information is instrumental in recognizing specific human activities.
Figure 3: The spectrogram of one subcarrier’s CSI amplitude for different activities: walking, picking up, and standing up (from left to right).
III-C2 Data denoising
CSI data is inherently replete with noise and often lacks the distinctness required to differentiate between various activities accurately. Consequently, the filtration of noise and extraction of salient features for classification through machine learning techniques is paramount. Prior research, including that by [45], has validated the effectiveness of PCA in this realm. PCA functions by reducing the dimensionality of extensive datasets, focusing primarily on key features that capture the bulk of signal information. Nevertheless, the efficacy of PCA is notably compromised in NLOS settings, where multipath fading and random noise are prevalent. Additionally, the CSI data exists in a high-dimensional vector space, characterized by its time series nature; it exhibits temporal variations ($T$ sampling packets) in conjunction with spatial variations ($N$ subcarriers). Extracting statistical information, particularly the temporal-spatial correlations, from these datasets presents a formidable challenge, one that conventional mathematical tools are ill-equipped to handle. To address this complexity, random matrix theory offers a mathematically rigorous solution for managing such data. In light of this, we introduce a novel denoising and feature extraction methodology: the high-dimensional factor model (HDFM), which is grounded in the Marčenko-Pastur (M-P) law [46].
The mathematical framework of HDFM is outlined as follows: Let $N$ represent the number of subcarriers, $T$ be the number of time sampling points, and $\mathbf{R}_{it}$ be the amplitude/phase of the $i$-th subcarrier at time $t$, where $i=1,2,\cdots,N$, $t=1,2,\cdots,T$, and $\frac{N}{T}\rightarrow c>0$. The CSI amplitude/phase matrix is expressed by Eq. (1), where $p$ is the number of temporal factors, $\mathbf{F}_{jt}$ denotes the $j$-th temporal factor at time $t$, and $\mathbf{L}_{ij}$ represents the loading of the $j$-th temporal factor on the $i$-th cross-sectional element. Furthermore, $\mathbf{U}_{it}$ stands for the additive random noise or idiosyncratic component.
$\mathbf{R}_{it}=\sum_{j=1}^{p}\mathbf{L}_{ij}\mathbf{F}_{jt}+\mathbf{U}_{it}.$ (1)
We construct the $p$-level residuals by removing temporal factors, using principal components:
$\mathbf{U}_{it}^{(p)}=\mathbf{R}_{it}-\mathbf{L}_{ij}^{(p)}\mathbf{F}_{jt}^{(p)},$ (2)
where $\mathbf{L}^{(p)}\mathbf{F}^{(p)}$ represents the estimated common factor obtained from $p$ principal components. The HDFM method lies in examining the eigenvalue distribution of the covariance matrix of $p$-level residuals $\mathbf{U}_{it}^{(p)}$:
$\mathbf{C}_{it}^{(p)}=\frac{1}{T}\mathbf{U}_{it}^{(p)}\mathbf{U}_{it}^{(p)H}.$ (3)
Upon accurately extracting temporal factors, the residual spectral density $\mathbf{C}_{it}^{(p)}$ closely aligns with the M-P law, and these factors are utilized as input to the classification algorithm. The pseudocode for the proposed method is delineated in Algorithm 1. For a comprehensive understanding, the proof of the HDFM is elaborated in Appendix B.
Algorithm 1 The HDFM method.
0:  CSI amplitude/phase data matrix: $\mathbf{R}_{it}$;
0:  CSI amplitude/phase features (temporal factors) $\mathbf{F}_{jt}^{(p)}$.
1:  Initialize the $p$-level residual matrix by Eq. (2);
2:  Compute the covariance matrix by Eq. (3);
3:  Calculate the eigenvalues $\mathbf{\lambda}$ of $\mathbf{C}_{it}^{(p)}$;
4:  Compute the eigenvalue distribution $\rho_{\mathbf{\lambda}}(p)$;
5:  Generate a Gaussian noise matrix $\hat{\mathbf{U}}_{it}\sim\mathcal{N}(0,\sigma^{2})$;
6:  Compute the covariance matrix for noise by Eq. (3);
7:  Compute the eigenvalues $\hat{\mathbf{\lambda}}$ of $\hat{\mathbf{C}}_{it}$;
8:  Compute the eigenvalue distribution $\rho_{\hat{\mathbf{\lambda}}}(\sigma)$;
9:  Minimize the distance between the two spectral distributions;
10:  Determine the number of temporal factors $p$ ;
11:  Compute the temporal factors $\mathbf{F}_{jt}^{(p)}$ as the final feature representation.
III-D Learning-based Algorithms
In our research, we integrate both machine learning (ML) and deep learning (DL) methodologies to evaluate the HDFM. We employ well-established ML algorithms like random forest (RF) [47], extreme gradient boosting (XGBoost) [48], and support vector machine (SVM) [49], alongside a novel dual-stream spatial-temporal attention deep learning framework designed for human activity classification. This framework’s architecture, as depicted in Fig. 4, begins with the preprocessing of CSI data through a spatial-temporal extractor. Following this, the extracted features are input into an attention layer, which utilizes an attention matrix to highlight the significance of various features and time steps. The integration of these features with the attention matrix elements results in a feature matrix with attention weights, which is then flattened into a feature vector. This vector facilitates the differentiation of action labels through a softmax classification layer.
Figure 4: Schematic of our proposed dual-stream spatial-temporal attention deep learning framework. Given CSI amplitude and phase features, we first use the spatial-temporal extractors to extract its activity-related features. Then, these features are fused using the attention mechanism and passed to a dense layer. Finally, the softmax is used to give output for the activity category prediction.
The attention mechanism model, initially developed for computer vision and based on the human ability to selectively concentrate on specific information, is adeptly applied in Wi-Fi-based passive sensing [50]. The high dimensionality of the features and the varying contributions of different features and time steps to the sensing task are managed through the incorporation of the attention mechanism, as substantiated by our time-frequency analysis illustrated in Fig. 3. The integration of this mechanism significantly enhances the accuracy of Wi-Fi sensing in our study.
In the architecture proposed, the network incorporates three convolutional layers followed by a LSTM layer, which serves to further refine the feature extraction process. As delineated in Table II, detailed specifications for each layer are provided. Specifically, the convolutional layers are configured with 32, 64, and 128 neurons, respectively, while the LSTM layer is equipped with 512 hidden units. The variable $M$ denotes the batch size, and $K$ signifies the number of labels.
TABLE II: Implementation details of the proposed network.
Layers Output dimension
Amplitude features Phase features
Input layer ($M$, 8, 5000) ($M$, 8, 5000)
Conv1D 1 ($M$, 32, 5000) ($M$, 32, 5000)
MaxPool1D 1 ($M$, 32, 2500) ($M$, 32, 2500)
Conv1D 2 ($M$, 64, 2500) ($M$, 64, 2500)
MaxPool1D 2 ($M$, 64, 1250) ($M$, 64, 1250)
Conv1D 3 ($M$, 128, 1250) ($M$, 128, 1250)
MaxPool1D 3 ($M$, 128, 625) ($M$, 128, 625)
LSTM ($M$, 625, 512) ($M$, 625, 512)
Concat ($M$, 625, 1024)
Attention layer ($M$, 1024)
Dense layer ($M$, $K$)
IV Experiments and Results
In this section, we implement and evaluate the performance of different methods as well as our proposed dual-stream spatial-temporal attention network model. All experiments are conducted on a server equipped with an Intel Xeon Gold 6138 CPU @ 2.00 GHz, 128 GB RAM, NVIDIA GeForce RTX 3090, operating on Linux and employing Python 3.8 and PyTorch 1.13.
Recognition performance can be measured from different perspectives using the following metrics: 1) Accuracy; 2) Precision; 3) Recall; and 4) F1-score. These metrics are defined as follows:
$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}},$ (4)
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}},$ (5)
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}},$ (6)
[…]
where TP, TN, FP, and FN denote true positive, true negative, false positive, and false negative, respectively. True positive signifies the accurate identification of positive instances by the model while true negative represents correctly identified negative instances. False positive denotes the count of negative instances erroneously classified as positive, and false negative indicates the count of positive instances incorrectly classified as negative by the model.
IV-A Performance Evaluation of RIS-enabled HAR
In this investigation, a comparative assessment is conducted to evaluate the efficacy of various algorithms within two distinct settings: environments without RIS configuration and those with RIS-optimized configurations, as elaborated in Table III. This examination highlights substantial enhancements in classifier accuracy following environmental optimization via RIS, with our proposed DS-STAN experiencing a significant accuracy uplift from 79.17% in non-configurable settings to 90.83% in RIS-optimized conditions. Comparable improvements are observed across other methodologies upon RIS activation, validating the critical role of RIS in augmenting the performance of activity-sensing technologies. Such advancements underscore the potential of RIS in revolutionizing wireless sensing systems, offering a pathway to achieve heightened precision and reliability in outcomes.
TABLE III: The overall recognition performances of different classification methods on non-configurable environment case and optimized-configurable environment case.
Methods RIS non-configurable environment RIS optimized-configurable environment
Accuracy Recall Precision F1-score Accuracy Recall Precision F1-score
RF-HDFM 0.5083 0.5148 0.5000 0.4837 0.8583 0.8649 0.8548 0.8547
XGBoost-HDFM 0.5292 0.5256 0.5265 0.5214 0.8417 0.8417 0.8364 0.8378
SVM-HDFM 0.5208 0.5200 0.5114 0.5116 0.8583 0.8598 0.8520 0.8546
DS-STAN-HDFM 0.7917 0.7907 0.7881 0.7859 0.9083 0.9084 0.9094 0.9089
IV-B Performance Evaluation of HDFM
In the preprocessing phase for CSI, PCA [17] and HDFM techniques are utilized for noise reduction and feature extraction. The extracted amplitude and phase features are subsequently input into various classifiers, including RF, XGBoost, SVM, and the proposed DS-STAN model that integrates time-frequency domain features. A comprehensive evaluation reveals HDFM’s significant enhancement of activity recognition accuracy in two datasets, with noticeable improvements over PCA-based approaches, as detailed in Table IV. Notably, the implementation of HDFM significantly enhances recognition accuracy in the Meeting room dataset, with improvements observed between 5.14% and 5.70% over PCA methods. Similarly, for the Office dataset, HDFM outperforms PCA, indicating accuracy gains ranging from 6.83% to 9.48%.
TABLE IV: Test accuracy comparison of different classification algorithms with two signal extraction methods on different datasets.
Datasets Methods Accuracy Recall Precision F1-score
Meeting room RF PCA 0.7736 0.7830 0.7837 0.7805
HDFM 0.8306 0.8375 0.8333 0.8350
XGBoost PCA 0.8014 0.8043 0.8133 0.8033
HDFM 0.8583 0.8654 0.8631 0.8634
SVM PCA 0.8236 0.8305 0.8283 0.8288
HDFM 0.8750 0.8782 0.8819 0.8797
DS-STAN PCA 0.9083 0.9128 0.9171 0.9110
HDFM 0.9639 0.9651 0.9636 0.9643
Office RF PCA 0.6948 0.7036 0.6867 0.6857
HDFM 0.7631 0.7714 0.7652 0.7643
XGBoost PCA 0.7219 0.7283 0.7120 0.7142
HDFM 0.8167 0.8218 0.8163 0.8181
SVM PCA 0.5646 0.5690 0.5561 0.5300
HDFM 0.6512 0.6508 0.6418 0.6278
DS-STAN PCA 0.9010 0.9011 0.8992 0.8991
HDFM 0.9726 0.9731 0.9738 0.9734
IV-C Performance Evaluation of Amplitude-phase Fusion
The comparative analysis reveals a significant performance boost in human activity recognition through the application of CSI amplitude-phase fusion data, as depicted in Fig. 5. This fusion approach outperforms amplitude-only and phase-only methods, demonstrating higher accuracy across different datasets. Notably, the Meeting room and Office datasets show accuracy of 96.39% and 97.26% respectively with the fusion method, affirming the hypothesis that combining amplitude and phase data substantially enhances recognition accuracy. This academic consolidation underlines the pivotal role of amplitude-phase data fusion in advancing CSI-based activity recognition systems.
(a) Performance comparisons of different methods on the Meeting room dataset.
(b) Performance comparisons of different methods on the Office dataset.
Figure 5: Performance comparisons of different methods on the two different datasets.
IV-D Performance Evaluation of Attention Mechanism
Fig. 6 reveals a pronounced improvement in activity recognition accuracy when utilizing a CSI amplitude-phase feature fusion method enhanced by an attention mechanism. Specifically, employing this mechanism increases accuracy to 96.39% and 97.26% for the Meeting room and Office datasets, respectively, a notable improvement from 90.14% and 90.12%. The attention mechanism’s impact is particularly pronounced in the Meeting room dataset, where the accuracy for the stand-up activity soared from 64% to 91%. These results highlight the critical advantage of integrating an attention layer, showcasing not only an increase in accuracy and stability but also the method’s adaptability and consistent performance across varied settings.
(a) Performance comparisons of with/without attention mechanism on the Meeting room dataset.
(b) Performance comparisons of with/without attention mechanism on the Office dataset.
Figure 6: Performance comparisons of attention mechanism on the two different datasets.
IV-E Performance Evaluation with Other Existing Methods
In our comparative analysis, presented in Table V, the performance of our proposed DS-STAN method for activity recognition, utilizing the Meeting room dataset, significantly outperforms existing benchmark classifiers, achieving an exceptional accuracy rate of 96.39%. This evidence not only affirms the superior efficacy of our approach but also highlights its potential to emerge as a forefront solution in the HAR domain.
TABLE V: Performance comparison of the proposed approach with other existing methods.
Authors Methods Accuracy
Yousefi et al. [18] LSTM 0.8708
Chen et al. [19] ABLSTM 0.7139
Feng et al. [21] Wi-multi 0.7514
Xiao et al. [22] DeepSeg 0.7389
Guo et al. [23] LCED 0.8681
Lu et al. [24] CeHAR 0.9000
Shalaby et al. [25] CNN-GRU-CNN 0.9194
Ours RISAR 0.9639
V Conclusion
Our research presents RISAR, an innovative device-free passive human activity recognition system leveraging Wi-Fi signals, tailored for NLOS environments. RISAR significantly enhances the accuracy of activity recognition. The system’s architecture adeptly integrates an HDFM for effective signal extraction and noise reduction, alongside an advanced dual-stream spatial-temporal attention network model. The experimental results from diverse environments demonstrate the system’s proficiency. Notably, RISAR shows remarkable improvements in recognition accuracy, evident in the transition from PCA-based methods to HDFM techniques. The introduction of an attention layer within the deep learning framework further augments this accuracy, underscoring the importance of feature selection and emphasis in complex signal environments.
In conclusion, RISAR represents a significant advancement in human activity recognition technology, offering enhanced accuracy and adaptability in various settings. Its ability to operate effectively in challenging NLOS scenarios and its adaptability to different environmental conditions mark it as a promising solution for future IoT applications in smart homes, healthcare, and beyond. This study not only contributes to the field of wireless sensing and activity recognition but also paves the way for future research exploring the integration of RIS and advanced machine learning techniques in other domains.
Appendix A Marčenko-Pastur law
Let $\mathbf{X}=\{x_{ij}\}$ be a $N\times T$ random matrix, whose entires are independent identically distributed (i.i.d.) variables with the mean $\mu(x)=0$ and the variance $\sigma^{2}(x)<\infty$. The corresponding covariance matrix is defined as $\mathbf{C}=\frac{1}{T}\mathbf{X}\mathbf{X}^{H}$. As $N,T\longrightarrow\infty$ but $c=\frac{N}{T}\in(0,1]$, according to M-P law, the empirical spectrum distribution of $\mathbf{C}$ converges to the limit with probability density function (PDF)
$f(x)=\begin{cases}\frac{1}{2\pi c\sigma^{2}x}\sqrt{(b-x)(x-a)},&a\leq x\leq b\\ 0,&\text{others}\end{cases}$ (8)
where $a=\sigma^{2}(1-\sqrt{c})^{2}$, $b=\sigma^{2}(1+\sqrt{c})^{2}$.
Appendix B Proof of the HDFM
Assumption 1
Let $\lambda_{1,T}\geq\ldots\geq\lambda_{p,T}$ represent the non-zero eigenvalues of $(\mathbf{L}\mathbf{F})(\mathbf{L}\mathbf{F})^{H}$. For $j=1,\ldots,p$, $\lambda_{j,T}\rightarrow\lambda_{j}$ as $T\rightarrow\infty$, where
$\lambda_{1}>\ldots>\lambda_{p}>\sigma^{2}\sqrt{c}.$
Theorem 1
Let $\hat{\lambda}_{1,T}\geq\ldots\geq\hat{\lambda}_{N,T}$ denote the eigenvalues of matrix $\mathbf{R}\mathbf{R}^{H}$. Under Assumption 1, for $j=1,\ldots,p$,
$\hat{\lambda}_{j,T}\xrightarrow[T\rightarrow\infty]{a.s.}\frac{(\lambda_{j}+\sigma^{2})(\lambda_{j}+\sigma^{2}c)}{\lambda_{j}}.$
Moreover,
$\hat{\lambda}_{p+1,T},\ldots,\hat{\lambda}_{N,T}\in[\sigma^{2}(1-\sqrt{c})^{2},\sigma^{2}(1+\sqrt{c})^{2}],$
almost surely (a.s.) for $T$ large enough.
Theorem 1 is derived from the general results established in [51]. It delineates a sudden transition in the behavior of the $j$-th dominant eigenvalue $\hat{\lambda}_{j,T}$ of $\frac{1}{T}\mathbf{R}\mathbf{R}^{H}$: if $\lambda_{j,T}\leq\sigma^{2}\sqrt{c}$, $\hat{\lambda}_{j,T}$ converges to the right-edge $\sigma^{2}(1+\sqrt{c})^{2}$ of the support defined by the Marčenko-Pastur law $\mu$, exhibiting no isolation. However, once $\lambda_{j,T}>\sigma^{2}\sqrt{c}$, $\hat{\lambda}_{j,T}$ converges to a limit outside the right edge of $\mu$, indicating isolation from the Marčenko-Pastur support.